OpenStudy (anonymous):

A large synchrotron at the famous Particle Accelerator in the deserts of Mexico has a diameter of 3.5 km. Protons that circle the synchrotron travel at the speed of light. a.) how long does it take the proton to complete one revolution b.) the protons enter the ring at an energy level of 10.0 GeV. They gain 1.5 MeV each revolution. c.) how long does it take the protons to be accelerated to 1000 GeV d.) how far do the protons travel dying this acceleration

OpenStudy (aaronq):

a) calculate the circumference, determine the time it would take to transit speed of light is 2.99 x10^8 m/s (though I'm pretty sure we can't reach the speed of light yet) b) is not a question c) if the enter with 10 GeV and gain 0.0015 GeV each revolution 1000 = 10 + 0.0015 x, x is number of revolutions d) use the circumference you got in part a) and in part d), multiply them

OpenStudy (anonymous):

so for part a.) just calculate the circumference?

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

k, i got 19.242 for the circumference :)

OpenStudy (aaronq):

you need units, always

OpenStudy (aaronq):

divide the circumference by the speed of light

OpenStudy (anonymous):

the speed of light is 2.99 x 10^8 m/s?

OpenStudy (anonymous):

what are the units?

OpenStudy (aaronq):

you're the one doing the calculation, you should know

OpenStudy (anonymous):

but is the speed of light 2.99 x 10^8 m/s? :P

OpenStudy (anonymous):

@Compassionate

OpenStudy (compassionate):

[One moment please...] [Processing question...]

OpenStudy (anonymous):

k btw i forgot to post up full question for part b.) its The protons enter the ring at an energy level of 10.0 GeV. They gain 1.5 MeV each revolution. How many revolutions must they travel before they reach 1000 GeV of energy?

OpenStudy (compassionate):

>How long does it take the proton to complete one revolution \[t = \frac{ d }{ v }\] \[t = \frac{ 3.5}{299,792.458 } \rightarrow 1.18\sec\] 10.0 GeV Gain 1.5 GeV each revolution >How many must they travel before they get 1,000 GeV \[1.5^{17} + 10 = 995.26GeV\] So it must make 17 revolutions. >How long does it take the protons to be accelerated to 1000 GeV They have to travel around 17 times, with a distance of 3.5km, at a speed of 299,792.458 km/s Using the time formula for time. 17 x 3.5km = 59.5km\[t = \frac{ 59.5 }{ 299,792.458 } \rightarrow 1.98\sec\] They travel 59.5 kilometers.

OpenStudy (anonymous):

thank you!! <33

OpenStudy (anonymous):

can you help me with another question?